Status Report

NASA Space Station On-Orbit Status 21 August 2004

By SpaceRef Editor
August 21, 2004
Filed under , , ,
NASA Space Station On-Orbit Status 21 August 2004

SpaceRef note: This NASA Headquarters internal status report, as presented here, contains additional, original material produced by (copyright © 2004) to enhance access to related status reports and NASA activities.

All ISS systems continue to function nominally, except those noted previously or below.  Another Saturday, with a light-duty crew schedule.

After breakfast at 2:40am EDT, CDR Padalka and FE Fincke performed the regular weekly 3-hr. station cleaning.  [“Uborka” includes removal of food waste products, cleaning of compartments with vacuum cleaner, wet cleaning of the Service Module (SM) dining table and other surfaces with “Fungistat” disinfectant and cleaning fan screens to avoid temperature rises.]

Immediately following, the crewmembers held their once-every-two-weeks teleconference with ISS Program Management (William Gerstenmaier) in Houston via S-band/audio.
Gennady took care of the daily routine inspection of the Service Module (SM)’s SOZh life support system (including SM/ASU toilet facility insert replacements).

Previous Reports

ISS On-orbit Status [HQ]
ISS Status [JSC]
Shuttle Processing [KSC]

In the U.S. Airlock, Mike Fincke terminated the recharging of the EMU batteries started yesterday.   [The recharge of batteries #2047 & #2048 in the BSA (battery stowage assembly) was done outside their regular 50-day maintenance cycle, in preparation for next week’s EMU/spacesuit repair activities.]

Mike also filled out the FFQ (food frequency questionnaire), which keeps an (almost-)regular weekly log of his nutritional intake over time on special MEC (medical equipment computer) software.

At 9:20 am, the crew held the WPC (weekly planning conference) with the ground, during which Mike and Gennady discussed next week’s “Look-Ahead Plan” (regularly prepared jointly by MCC-H and MCC-M planners), via S-band/audio.

Padalka was kept busy with several optional tasks waiting on the Russian task list for today, leading off with his fourth operations and measurement session of the Molniya-SM/LSO hardware, already installed at SM window #3.  [Objective of the often-repeated geophysical Molniya-SM, similar to the French LSO experiment, is to record storm phenomena and other related events in the Earth’s equatorial regions.  The payload uses the French-provided EGE-1 laptop running the latest NORAD orbital parameters (TLEs, two-line elements).  Once Gennady started the recording session, the payload works automatically until 8/24.  Its data will then be copied to HDD (hard disk drive) on 8/29.]

The second task for the CDR was another session of the “Diatomeya” ocean observations program, for which he used the DSR PD-150P video camera and Nikon F5 digital still camera with 24/85-mm lens from SM windows #7 and #8 to collect photo and video data on bioproductive regions (algae blooms) that appear in the last ten days of August in the North Atlantic areas.  [Today’s “must” target was the North Atlantic’s Gulf Stream delta.]

Gennady also performed a run of the Russian “Ekon” earth photography program, today taking Nikon D1 (800 mm lens) pictures of the city of Sevastopol.

A fourth item on Padalka’s task list was photographic documentation of the Matryoshka-R radiation experiment’s “Phantom” unit and SPD passive sensors in the Russian segment, which he deployed yesterday.

Another task-listed job for Gennady was a session with the VC6 “Delta” program’s ETD experiment (Investigation of the Coordination of Eye and Head Movements).   [After a calibration with the calibrating unit, the experiment investigates horizontal eye and head movement coordination, measured Listing’s plane, and determined the orientation of the vestibulo-ocular coordinate system, using five target marks on the horizontal plane.  Each step required another prior calibration run, using visual target cues or the calibration unit.]

The CDR also performed another task-listed session with the Uragan (“hurricane”) earth-imaging program, using the Kodak 760 DSC (digital still camera) with 800mm-lens from SM windows.  [Today’s targets for his photo imagery were the Turkish Black Sea coast with the cities of Sinop, Samsun, Ordu and Trabzon, the Red Sea coast with islands, the city of Lagos, Nigeria, and the Huascaran volcano in Peru.]

For his chosen Saturday Science program today, FE/SO Mike Fincke conducted a checkout of the BSTC (Biotechnology Specimen Temperature Controller).   [In the procedures, he connected hoses/cables, then turned the gas supply on and watched the BSTC temperature rise to its set point.  Specialists monitored the hardware on the ground to ensure that everything was functioning properly, in preparation for experiments in future Increments.]

Mike Fincke’s attempt yesterday to check out the new RSP (Respiratory Support Pack) #1004 that arrived on 15P ran into a problem, which engineers are investigating.   [RSP #1004 is currently considered failed and may not be used for any medical contingency activities.  The old RSP #1003, still in the CheCS (crew health care systems) rack, is to be used for any medical contingencies until further notice.]

A new Russian evaluation of acoustic measurements made by the crew in the last few weeks points out the need for lowering the noise level in the RS and recommends that crewmembers should continue using protective devices when working in places where noisy equipment is located as well as during sleep periods (especially when doors to sleep compartments [“kayutas”] are left open).

Update on Node smoke detector #2:  SD-2 in the Node remains inhibited (SD-1 is working OK).  Experts agree that the most likely explanation for its spurious behavior, which caused a false fire alarm on 8/18, is due to dust or debris that has entered the housing. This can result in elevated readings of scattered particles, which is what has been observed on that detector.  The detector’s readings have returned to their normal values, and it could be re-enabled right now, but engineers, in the name of conservatism, will watch the readings some more to make sure no other scatter spikes occur.  The plan is to re-enable SD-2 early next week.

Weekly Science Update (Expedition Nine — 17th):

GASMAP:  Nothing new.

Human Research Facility/Workstation (HRF WS):  Continuing.

Advanced Ultrasound (ADUM):  Nothing new.

Hand Posture Analyzer (HPA:  Nothing new.

In-Space Soldering Investigation (ISSE):  Nothing new.

Pore Formation and Mobility Investigation (PFMI):  Nothing new.
Space Acceleration Measurement System (SAMS):  SAMS captured the Progress 15 docking and reboost.
Microgravity Acceleration Measurement System (MAMS):  MAMS captured quasi-steady data during the Progress 15 docking and reboost.

Protein Crystal Growth-Single Locker Thermal Enclosure System (PCG-STES):  Nominal.

Protein Crystal Growth Monitoring by Digital Holographic Microscope (PromISS):  Nothing new.

Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions (InSPACE):    Planned.

Binary Colloidal Alloy Test-3 (BCAT-3):   Mike was thanked again for his continued support of BCAT-3 and for taking the time to gather additional data.  “The images have begun to be received by the PIs.   Please hold on to the PCMCIA card until archival is completed.”

Renal Stone (RS):  Nothing new.

Synchronized Position Hold, Engage, Reorient, Experimental Satellites (SHERES):  Nothing new.

Foot/Ground Reaction Forces during Space Flight (FOOT):  Nothing new.

Materials ISS Experiment (MISSE):  In progress. Deployed outside on the U.S. Airlock.  Nominal and collecting data.

Cellular Biotechnology Support Systems-Fluid Dynamics Investigation (CBOSS-FDI):  Nothing new.

Biotechnology Specimen Temperature Controller (BSTC):  Mike was thanked for choosing BSTC Check-out for today’s Saturday Science.  “Because of your extra help, our hardware will be ready for more science on future Increments”.

Yeast Group Activation Packs (Yeast GAP):  Nothing new.

Earth Knowledge Acquired by Middle School Students (EarthKAM):  Nothing new.

Earth Science Toward Exploration Research (ESTER):  Nothing new.

Serial Network Flow Monitor (SNFM):  Nothing new.

Fluid Merging Viscosity Measurement (FMVM):  Nothing new.

Viscous Liquid Foam–Bulk Metallic Glass (Foam):   Nothing new.

BIOPSY (Effects of Prolonged Space Flight on Human Skeletal Muscle):  Nothing new.

Coarsening in Solid-Liquid Mixtures-2 (CSLM-2):  Planned.

Commercial Generic Bioprocessing Apparatus (CGBA):  Nothing new.

Miscible Fluids in Microgravity (MFMG):  Nothing new.

Educational Payload Operations (EPO):  “The Chicken Shake demonstration was excellent.  Maryland Science Center looks forward to including it in educator workshops and student programs.  We appreciate your enthusiastic support of education.”

Capillary Flow Experiment (CFE):    The CFE team is  developing new experiments based on the 8/17 discussion with Mike of possible extra science on future Saturdays.

Crew Earth Observations (CEO):  A mosaic of two of ISS/CEO images from last month of China’s Yellow River delta is being published on NASA’s Earth Observatory website this week.  These images are used to illustrate the intense human development of this region for agriculture, commerce, and petroleum production as well as the dynamics of the river and its sediment and the strikingly colored plankton bloom occurring just offshore at the time of your pass.  Despite the station’s adverse attitude, the crew has continued to take many useful and interesting photos.  Their long-lens views of Assam have been cataloged and are available on the web page (see below).  The CEO team is hopeful with improving weather in the coming weeks to find better opportunities to take less oblique photos of the Assam region for your family.

Today’s optional CEO photo targets, in the current LVLH attitude no longer limited by Flight Rule constraints on the use of the Lab nadir/science window, except for the shutter closure and condensation-prevention plan (limited to 90 min. in 24 hours) were Lahore, Pakistan (ISS had a nadir pass over the Lahore metropolitan area, and weather should have been clear. This was a good chance for acquisition of data for land cover/land use change analysis), Internal waves, Azores Is – Mid Atlantic (weather was clear over the mid-Atlantic and should have afforded a good opportunity for internal waves.  Looking left of track, and the sunglint point was almost directly beneath the station), and Internal waves, Western Equatorial Atlantic (this overpass presented a good opportunity to capture internal waves off the NE coast of South America.  Looking right of track, with the sunglint point slightly behind the ISS).

CEO images can be viewed at these websites:

See also the website “Space Station Challenge” at:

To view the latest photos taken by the expedition 9 crew visit:

Expedition 9 Flight Crew Plans can be found at

Major upcoming events:

  • ISS Reboost-2 — 8/24 (delta-V = 2.2 m/s);
  • EVA-11 — 9/3;
  • Soyuz 9S launch — 10/9;
  • Soyuz 9S dock — 10/11;
  • Soyuz 8S undock/land — 10/19;
  • Soyuz 9S relocate — 11/18;
  • Progress 16P launch — 11/24.

Previous NASA ISS On-orbit Status Reports can be found here. Previous NASA Space Station Status Reports can be found here. Previous NASA Space Shuttle Processing Status Reports can be found here. A collection of all of these reports and other materials relating to Return to Flight for the Space Shuttle fleet can be found here.

ISS Location NOW

Full Size/Update
Real Time ISS TrackerMore Links

ISS Orbit  (as of this morning, 8:36am EDT [= epoch]):

  • Mean altitude — 357.9 km
  • Apogee height — 363.2 km
  • Perigee height — 352.7 km
  • Period — 91.7 min.
  • Inclination (to Equator) — 51.63 deg
  • Eccentricity — 0.0007791
  • Solar Beta Angle — -9.8 deg (magnitude decreasing)
  • Orbits per 24-hr. day — 15.7
  • Mean altitude loss in last 24 hours — 150 m
  • Revolutions since FGB/Zarya launch (Nov. ’98)  — 32858

ISS Altitude History

Apogee height Mean AltitudePerigee height

ISS Altitude History

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see In addition, information on International Space Station sighting opportunities can be found at on NASA’s Human Spaceflight website. The current location of the International Space Station can be found at at NASA’s Marshall Space Flight Center. Additional satellite tracking resources can be found at

SpaceRef staff editor.